# Large Language Model Inference Acceleration: A Comprehensive Hardware Perspective

### v3

@@ 3.4.3. ASIC
Fusion. In 2020, Groq company introduces the Tensor Streaming Processor (TSP) architecture (abts2020think, ), a functionally-sliced microarchitecture with memory units interleaved with vector and matrix compute units in order to take advantage of dataflow locality. The first TSP implementation yields a computational density of more than 1 TOPS/$mm^{2}$ for its 25$\times$29 mm 14nm chip at 900 MHz. In 2022, Groq company introduces a novel software-defined communication approach for large-scale interconnection networks of TSP elements (abts2022software, ). This scalable communication fabric is based on a software-defined dragonfly topology, ultimately yielding a parallel machine learning system with elasticity to support a variety of workloads. Each TSP contributes 220 MB to the global memory capacity, with the maximum capacity limited only by the network’s scale. The large-scale parallel system achieves up to 10,440 TSPs and more than 2 TB of global memory accessible in less than 3ms of end-to-end system latency. Based on two previous works, Groq Language Processing Unit (LPU) (groq_lpu, ) is fabricated in 14nm with 185W power consumption. According to (groq_lpu_result, ), LPU achieves 814 tokens/s for Gemma-7B model. Another commercial company, HyperAccel, also proposes a LLM inference chip with dataflow architecture. Latency Processing Unit (LPU) (moon2024lpu, ) introduces streamlined hardware that maximizes the effective memory bandwidth usage during end-to-end inference regardless of the model size to achieve up to 90% bandwidth utilization for high-speed text generation. It consists of expandable synchronization link (ESL) that hides bulk of the data synchronization latency in a multi-device system to achieve nearperfect scalability. Its on-chip power is 0.284W sythesised by Samsung 4nm and the total system power is 86W with 96GB HBM3. For OPT-1.3B/6.7B/13B/30B, LPU achieves 769.23 tokens/s, 217.39 tokens/s, 112.40 tokens/s and 49.26 tokens/s, respectively. The Wafer-Scale Engine (WSE-3) (lie2024wafer, ), which powers the Cerebras CS-3 system, is the largest chip ever built. The WSE-3 is 57$\times$ larger than the largest GPU, has 52$\times$ more compute cores, and 880$\times$ more high performance on-chip memory. The only wafer scale processor ever produced, it contains 4 trillion transistors, 900,000 AI-optimized cores, and 44 gigabytes of high performance on-wafer memory. Each wafer consists of 84 dies with 40GB on-chip memory and 97W power consumption. For Llama3.1-8B, it can achieve about 1800 tokens/s, which is 20$\times$ faster than hyperscale clouds.
@@
Coarse-grained Processing. Tensor Contraction Processor (TCP) (kim2024tcp, ), is composed of coarse-grained PEs, which are designed to be flexible enough to be configured as a large-scale single unit or a set of small independent compute units. TCP chip is designed and fabricated in 5nm technology with 256MB SRAM and 1.5 TB/s 48GB HBM3 under 150W. For Llama2-7B model, TCP achieves about 125 tokens/s with batch size 1 and 1176 tokens/s with batch size 8. Habana Gaudi (habana_gaudi, ) and Gaudi2 (habana_gaudi2, ) consists of two main compute engines, Matrix Multiplication Engine (MME) and Tensor Processor Core (TPC) cluster. The TPC unit is a SIMD processor tailor-made for general deep learning operations. The biggest difference between GPU and Gaudi architecture is the size of MME. Gaudi architecture can handle 256$\times$256 matrix multiplication, which requires 512 input elements per cycle while GPU architecture with 16$\times$16 Tensor Core requires 8K input elements. Therefore, Gaudi architecture can save 16$\times$ less read bandwidth requirements. Besides compute engine, Gaudi2 includes 96 GB of HBM2E memories delivering 2.45 TB/sec bandwidth, in addition to a 48 MB of local SRAM. Gaudi is fabricated in 16nm and Gaudi2 is fabricated in 7nm technique node with 600W power consumption. For Llama2-7B and Llama2-13B models, Gaudi2 achieves from 81.97 to 111.11 tokens/s and from 49.02 to 64.52 tokens/s with batch size 1, respectively.
3.4.4. PIM/NDP
@@
Nonlinear Function Approximation. AttentionLego (cong2024attentionlego, ) proposes a PIM-based matrix-vector multiplication and look-up table-based Softmax design. PIM-GPT (wu2024pim, ), which achieves end-to-end acceleration of GPT inference with high performance and high energy efficiency. At the hardware level, PIM-GPT is a hybrid system that includes DRAM-based PIM chips to accelerate VMM near data and an application-specific integrated circuit (ASIC) to support other functions that are too expensive for PIM including necessary nonlinear functions, data communication, and instructions for the PIM chips. At the software level, the mapping scheme is optimized to efficiently support the GPT dataflow. Overall, PIM-GPT achieves 89$\times$ speedup and 221$\times$ energy efficiency over NVIDIA T4 GPU on 8 GPT models with up to 1.4 billion parameters. SAL-PIM (han2024sal, ) also adopts LUT-based linear interpolation to perform complex nonlinear functions in PIM.
Figure 10. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with operator optimization.
Storage Optimization. By observing that a key impediment to truly harness PIM acceleration is deducing optimal data-placement to place the matrix in memory banks, PIMnast (ibrahim2024balanced, ) proposes matrix tiling/ordering algorithms to tackle these factors and identify orchestration knobs that impact PIM acceleration. For OPT-6.7B model, compared to the SoC baseline (AMD Ryzen PRO 7040 Series processors comprising eight CPU cores, 12 compute units of GPU cores, 16 AIE tiles, and eight channels of LPDDR5 memory for a peak memory bandwidth of 120 GB/s), PIMnast achieves 4.5$\times$ speedup for per-token latency.
3.4.5. Comparison
In Figure 10, for operator optimization, power consumption ranges from 10.44W to 600W, with inference speeds between 4.13 tokens/s and 1800 tokens/s. The energy efficiency ranges from 0.092 tokens/J to 18.556 tokens/J. MARCA (li2024marca, ) (ASIC) achieves the lowest power consumption with 10.44W. Cerebras WSE-3 (lie2024wafer, ) (ASIC) achieves the highest throughput with batch size 1 (post-silicon results) and the highest energy efficiency.
  • For GPUs, power consumption ranges from 32W to 400W, with inference speeds between 4.13 tokens/s and 145.04 tokens/s. The energy efficiency ranges from 0.092 tokens/J to 0.522 tokens/J.
  • For ASICs, power consumption ranges from 10.44W to 600W, with inference speeds between 11.89 tokens/s and 1800 tokens/s, found in the upper left section of the figure. The energy efficiency ranges from 0.108 tokens/J to over 18.556 tokens/J, outperforming the GPU platform.
3.5. Heterogeneous Cooperation
3.5.1. Overview
@@
Figure 11. Heterogeneous cooperation.
@@
Memory Enhancement. focuses on two key aspects: storage capacity enhancement and bandwidth enhancement. As we move into the LLM eras, computational tasks are becoming increasingly complex, leading to greater demands on memory capacity and bandwidth. From the perspective of storage capacity, memory enhancement is achieved by integrating more on-chip memory by using 3D storage stacks on the same chip area. This allows more data and models to be totally stored in a limited on-chip space, reducing the need to access external storage and improving overall efficiency. Another memory enhancement is to accelerate memory-bound computations by increasing data transfer speeds. For heterogeneous cooperation, bandwidth often restricts computational performance. To address this issue, much methods are proposed to support high-speed on-chip (li20204gbps, ; li2023method, ; li2021gain, ) and chip-to-chip interconnect, such as Compute Express Link (CXL) (sharma2022compute, ) and NVLink (foley2017ultra, ). These techniques enable low-latency, high-bandwidth data transmission between different chips, significantly speeding up data exchange processes.
Table 7. Heterogeneous cooperation on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@ 3.5.2. CPU
Computing Enhancement. (1) CPU-GPU. Only using the CPU may result in slower performance, so many methods employ a combination of CPU and GPU to enhance LLM inference speed. For personal computers, PowerInfer (song2023powerinfer, ) proposes that the hot-activated neurons should be preloaded onto the GPU for fast access, while cold-activated neurons are computed on the CPU, thus significantly reducing GPU memory demands and CPU-GPU data transfers. PowerInfer further integrates adaptive predictors and neuron-aware sparse operators. For various models (OPT-30B/66B, Falcon-40B, and Llama-70B) on a Intel i9-13900K processor equipped with an NVIDIA RTX 4090, it achieves 8.32 tokens/s on average. On a Intel i7-12700K processor equipped with an NVIDIA RTX 2080Ti, it achieves 5.77 tokens/s on average. For smartphones, PowerInfer-2 (xue2024powerinfer, ) further leverages the highly heterogeneous XPUs present in smartphone SoCs, such as asymmetric big.LITTLE CPU cores, GPU, and NPU. On OnePlus 12 equipped with SnapDragon 8 Gen 3, PowerInfer-2 achieves 11.7 tokens/s and 10.5 tokens/s on Llama-7B and Mistral-7B, respectively. On OnePlus Ace 2 equipped with SnapDragon 8+ Gen 1, PowerInfer-2 achieves 6.5 tokens/s and 6.3 tokens/s on Llama-7B and Mistral-7B, respectively. For servers, Kim et al. (kim2024exploiting, ) propose an adaptive model to determine the LLM layers to be run on CPU and GPU, respectively, based on the memory capacity requirement and arithmetic intensity. They then propose CPU-GPU cooperative computing that exploits the AMX extensions of the latest Intel CPU. The evaluation demonstrates that for OPT-30B model, CPU-GPU cooperative computing achieves 336 tokens/s with batch size 1 and input tokens 2016 in prefill stage. In decode stage, it achieves about 25 tokens/s with batch size 90 and input tokens 2016, and about 300 tokens/s with batch size 1400 and input tokens 64.
3.5.3. PIM/NDP
@@
Computing Enhancement. (1) PIM-NPU. NeuPIMs (heo2024neupims, ) and IANUS (seo2024ianus, ) are heterogeneous PIM acceleration systems that jointly exploits a conventional GEMM-focused NPU and GEMV-optimized PIM devices. NeuPIMs (heo2024neupims, ) first proposes dual row buffers in each bank, facilitating the simultaneous management of memory read/write operations and PIM commands, to enable concurrent operations on both NPU and PIM platforms. Further, it employs a runtime sub-batch interleaving technique to maximize concurrent execution for the inherent dependencies between GEMM and GEMV in LLMs. Our evaluation demonstrates that NeuPIMs with $\sim$76W (0.6348W+75W) power consumption (32GB HBM) achieves $\sim$3k tokens/s with batch size 8 for GPT3-7B. IANUS (seo2024ianus, ) proposes novel PIM access scheduling that manages not only the scheduling of normal memory accesses and PIM computations but also workload mapping across the PIM and the NPU. The evaluations show that for GPT-6.7B model, IANUS achieves 127.1 tokens/s with about 240W power consumption. Cambricon-LLM (yu2024cambricon, ) proposes a chiplet-based hybrid architecture with NPU and a dedicated NAND flash chip to enable efficient LLM inference. It utilizes both the high computing capability of NPU and the data capacity of the NAND flash chip. The NAND flash chip is enhanced by in-flash computing to perform precise lightweight on-die processing, and the NPU performs matrix operations and handles special function computations. Experimental results show that Cambricon-LLM with $\sim$37W power consumption achieves 3.44 tokens/s ($\sim$) and 36.34 tokens/s for 70B and 7B LLMs, which is over 22$\times$ to 45$\times$ faster than existing flash-offloading technologies, respectively. (2) PIM-GPU. MoNDE (kim2024monde, ) is a near-data computing solution that efficiently enables Mixture-of-Experts (MoE) LLM inference on heterogeneous hardwares. MoNDE reduces the volume of MoE parameter movement by transferring only the hot experts to the GPU, while computing the remaining cold experts inside the host memory device. MoNDE can achieve inference latency comparable to an ideal GPU system with infinite memory. AttAcc (choi2023unleashing, ; park2024attacc, ) is also an heterogeneous system equipped with GPUs to accelerate the attention layers. Compared to the monolithic state-of-the-art GPU system, AttAcc achieves significantly higher throughput (up to 2.81$\times$) and energy efficiency (up to 2.67$\times$) for GPT3-175B model. In 2024, SK Hynix proposes LPDDR6-based heterogeneous LLM system called AiMX-xPU(kim2024sk, ), which consists 1 NVIDIA H100 GPU and 3 AiMX, achieving 167 tokens/s, 220 tokens/s, and 900 tokens/s with batch size 1, 8, and 32 for OPT-30B. (3) PIM-FPGA. Kang et al. (kang2023era, ) put the memory-bound GEMV calculations including projections, feed-forward-network layer, and the last fully-connected layer with pre-trained weight on HBM2-PIM and evaluate system with PIM-powered Xilinx Alveo U280 board. For GPT-1.3B model, after scaling the bandwidth by 8$\times$ for PIM technology, it can achieve about 347.83 tokens/s, which is 1.6$\times$ faster than the NVIDIA A100 GPU. From 2022 to 2023, SK Hynix proposes GDDR6-based accelerator-in-memory named AiM (lee20221ynm, ; kwon2022system, ) and integrates it with 2 XLINX XCVU9P FPGA chips for control and communication as a prototype called AiMX (AiMX, ) to achieve 330 tokens/s for OPT-6.7B with batch size 1. (4) PIM-TPU. H3D-Transformer (luo2024h3d, ) proposes a heterogeneous 3D-based accelerator design for transformer models, which adopts an interposer substrate with multiple 3D memory/logic hybrid cubes optimized for accelerating different MatMul workloads. An approximate computing scheme is proposed to take advantage of heterogeneous computing paradigms of mixed-signal compute-in-memory (CIM) and digital tensor processing units (TPU). From the system-level evaluation results, 10 TOPS/W energy efficiency is achieved for the GPT2 model, which is about 2.6$\times$-3.1$\times$ higher than the baseline with 7nm TPU and stacked FeFET memory. 3D-HI (sharma2023heterogeneous, ) leverage chiplet-based heterogeneous integration to design a Network-on-Interposer (NoI) architecture to accelerate LLM inference. 3D-HI uses the streaming multiprocessors along with the associated memory controllers (SM-MCs) for multi-head attention and ReRAM cores for the feed-forward network, which optimize both achievable energy efficiency and throughput. Further, vertical integration on top of a 2.5D interposer helps to enhance overall system performance and alleviates the issue of memory bottlenecks. Experimental results demonstrate that 3D-HI lowers the latency and energy consumption by up to 22.8$\times$ and 5.36$\times$ with respect to an equivalent state-of-the-art chiplet-based platform.
@@
Figure 12. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with heterogeneous cooperation.

### v4

@@ 3.4.5 ASIC
Fusion. In 2020, Groq company introduces the Tensor Streaming Processor (TSP) architecture [294], a functionally-sliced microarchitecture with memory units interleaved with vector and matrix compute units in order to take advantage of dataflow locality. The first TSP implementation yields a computational density of more than 1 TOPS/$mm^{2}$ for its 25$\times$29 mm 14nm chip at 900 MHz. In 2022, Groq company introduces a novel software-defined communication approach for large-scale interconnection networks of TSP elements [295]. This scalable communication fabric is based on a software-defined dragonfly topology, ultimately yielding a parallel machine learning system with elasticity to support a variety of workloads. Each TSP contributes 220 MB to the global memory capacity, with the maximum capacity limited only by the network’s scale. The large-scale parallel system achieves up to 10,440 TSPs and more than 2 TB of global memory accessible in less than 3ms of end-to-end system latency. Based on two previous works, Groq Language Processing Unit (LPU) [76] is fabricated in 14nm with 185W power consumption. According to [296], LPU achieves 814 tokens/s for Gemma-7B model. Another commercial company, HyperAccel, also proposes a LLM inference chip with dataflow architecture. Latency Processing Unit (LPU) [244] introduces streamlined hardware that maximizes the effective memory bandwidth usage during end-to-end inference regardless of the model size to achieve up to 90% bandwidth utilization for high-speed text generation. It consists of expandable synchronization link (ESL) that hides bulk of the data synchronization latency in a multi-device system to achieve nearperfect scalability. Its on-chip power is 0.284W sythesised by Samsung 4nm and the total system power is 86W with 96GB HBM3. For OPT-1.3B/6.7B/13B/30B, LPU achieves 769.23 tokens/s, 217.39 tokens/s, 112.40 tokens/s and 49.26 tokens/s, respectively. Based on Cerebras WSE2, WaferLLM [252] is the first wafer-scale LLM inference system. For full LLM inference, WaferLLM achieves 2700 tokens/s and 2039 tokens/s on Llama3-8B and Llama2-13B with 434W and 354W, respectively. The Wafer-Scale Engine (WSE-3) [77], which powers the Cerebras CS-3 system, is the largest chip ever built. The WSE-3 is 57$\times$ larger than the largest GPU, has 52$\times$ more compute cores, and 880$\times$ more high performance on-chip memory. The only wafer scale processor ever produced, it contains 4 trillion transistors, 900,000 AI-optimized cores, and 44 gigabytes of high performance on-wafer memory. Each wafer consists of 84 dies with 40GB on-chip memory and 15kW power consumption. For Llama3.1-8B, it can achieve about 1800 tokens/s, which is 20$\times$ faster than hyperscale clouds.
@@
Coarse-grained Processing. Tensor Contraction Processor (TCP) [247], is composed of coarse-grained PEs, which are designed to be flexible enough to be configured as a large-scale single unit or a set of small independent compute units. TCP chip is designed and fabricated in 5nm technology with 256MB SRAM and 1.5 TB/s 48GB HBM3 under 150W. For Llama2-7B model, TCP achieves about 125 tokens/s with batch size 1 and 1176 tokens/s with batch size 8. Habana Gaudi [248] and Gaudi2 [249] consists of two main compute engines, Matrix Multiplication Engine (MME) and Tensor Processor Core (TPC) cluster. The TPC unit is a SIMD processor tailor-made for general deep learning operations. The biggest difference between GPU and Gaudi architecture is the size of MME. Gaudi architecture can handle 256$\times$256 matrix multiplication, which requires 512 input elements per cycle while GPU architecture with 16$\times$16 Tensor Core requires 8K input elements. Therefore, Gaudi architecture can save 16$\times$ less read bandwidth requirements. Besides compute engine, Gaudi2 includes 96 GB of HBM2E memories delivering 2.45 TB/sec bandwidth, in addition to a 48 MB of local SRAM. Gaudi is fabricated in 16nm and Gaudi2 is fabricated in 7nm technique node with 600W power consumption. For Llama2-7B and Llama2-13B models, Gaudi2 achieves from 81.97 to 111.11 tokens/s and from 49.02 to 64.52 tokens/s with batch size 1, respectively.
3.4.6 PIM/NDP
@@
Nonlinear Function Approximation. AttentionLego [254] proposes a PIM-based matrix-vector multiplication and look-up table-based Softmax design. PIM-GPT [255], which achieves end-to-end acceleration of GPT inference with high performance and high energy efficiency. At the hardware level, PIM-GPT is a hybrid system that includes DRAM-based PIM chips to accelerate VMM near data and an application-specific integrated circuit (ASIC) to support other functions that are too expensive for PIM including necessary nonlinear functions, data communication, and instructions for the PIM chips. At the software level, the mapping scheme is optimized to efficiently support the GPT dataflow. Overall, PIM-GPT achieves 89$\times$ speedup and 221$\times$ energy efficiency over NVIDIA T4 GPU on 8 GPT models with up to 1.4 billion parameters. SAL-PIM [256] also adopts LUT-based linear interpolation to perform complex nonlinear functions in PIM. SoftmAP [166] deploys integer only low-precision Softmax on In-Memory Compute hardware to accelerate integer-only LLM inference. It achieves up to 1000$\times$ improvement in the energy-delay product compared to A100 and RTX3090 GPUs.
Figure 10: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with operator optimization.
Storage Optimization. By observing that a key impediment to truly harness PIM acceleration is deducing optimal data-placement to place the matrix in memory banks, PIMnast [253] proposes matrix tiling/ordering algorithms to tackle these factors and identify orchestration knobs that impact PIM acceleration. For OPT-6.7B model, compared to the SoC baseline (AMD Ryzen PRO 7040 Series processors comprising eight CPU cores, 12 compute units of GPU cores, 16 AIE tiles, and eight channels of LPDDR5 memory for a peak memory bandwidth of 120 GB/s), PIMnast achieves 4.5$\times$ speedup for per-token latency.
3.4.7 Comparison
In Figure 10, for operator optimization, power consumption ranges from 10.44W to 700W, with inference speeds between 4.13 tokens/s and 2700 tokens/s. The energy efficiency ranges from 0.034 tokens/J to 12.34 tokens/J. MARCA [246] (ASIC) achieves the lowest power consumption with 10.44W. WaferLLM [252] (ASIC) achieves the highest throughput with batch size 1 (post-silicon results) and the highest energy efficiency.
  • For CPUs, power consumption ranges from 120W to 350W, with inference speeds between 4.32 tokens/s and 12 tokens/s. The energy efficiency ranges from 0.034 tokens/J to 0.036 tokens/J.
  • For GPUs, power consumption ranges from 32W to 700W, with inference speeds between 4.13 tokens/s and 181 tokens/s. The energy efficiency ranges from 0.092 tokens/J to 0.522 tokens/J.
  • For ASICs, power consumption ranges from 10.44W to 600W, with inference speeds between 23.78 tokens/s and 2700 tokens/s, found in the upper left section of the figure. The energy efficiency ranges from 0.108 tokens/J to over 12.34 tokens/J, outperforming the GPU platform.
3.5 Heterogeneous Cooperation
3.5.1 Overview
@@
Figure 11: Heterogeneous cooperation.
@@
Memory Enhancement. focuses on two key aspects: storage capacity enhancement and bandwidth enhancement. As we move into the LLM eras, computational tasks are becoming increasingly complex, leading to greater demands on memory capacity and bandwidth. From the perspective of storage capacity, memory enhancement is achieved by integrating more on-chip memory by using 3D storage stacks on the same chip area. This allows more data and models to be totally stored in a limited on-chip space, reducing the need to access external storage and improving overall efficiency. Another memory enhancement is to accelerate memory-bound computations by increasing data transfer speeds. For heterogeneous cooperation, bandwidth often restricts computational performance. To address this issue, much methods are proposed to support high-speed on-chip [299, 300, 301] and chip-to-chip interconnect, such as Compute Express Link (CXL) [302] and NVLink [62]. These techniques enable low-latency, high-bandwidth data transmission between different chips, significantly speeding up data exchange processes.
Table 7: Heterogeneous cooperation on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@ 3.5.2 CPU
Computing Enhancement. (1) CPU-GPU. Only using the CPU may result in slower performance, so many methods employ a combination of CPU and GPU to enhance LLM inference speed. For personal computers, PowerInfer [259] proposes that the hot-activated neurons should be preloaded onto the GPU for fast access, while cold-activated neurons are computed on the CPU, thus significantly reducing GPU memory demands and CPU-GPU data transfers. PowerInfer further integrates adaptive predictors and neuron-aware sparse operators. For various models (OPT-30B/66B, Falcon-40B, and Llama-70B) on a Intel i9-13900K processor equipped with an NVIDIA RTX 4090, it achieves 8.32 tokens/s on average. On a Intel i7-12700K processor equipped with an NVIDIA RTX 2080Ti, it achieves 5.77 tokens/s on average. For smartphones, PowerInfer-2 [260] further leverages the highly heterogeneous XPUs present in smartphone SoCs, such as asymmetric big.LITTLE CPU cores, GPU, and NPU. On OnePlus 12 equipped with SnapDragon 8 Gen 3, PowerInfer-2 achieves 11.7 tokens/s and 10.5 tokens/s on Llama-7B and Mistral-7B, respectively. On OnePlus Ace 2 equipped with SnapDragon 8+ Gen 1, PowerInfer-2 achieves 6.5 tokens/s and 6.3 tokens/s on Llama-7B and Mistral-7B, respectively. For servers, Kim et al. [258] propose an adaptive model to determine the LLM layers to be run on CPU and GPU, respectively, based on the memory capacity requirement and arithmetic intensity. They then propose CPU-GPU cooperative computing that exploits the AMX extensions of the latest Intel CPU. The evaluation demonstrates that for OPT-30B model, CPU-GPU cooperative computing achieves 336 tokens/s with batch size 1 and input tokens 2016 in prefill stage. In decode stage, it achieves about 25 tokens/s with batch size 90 and input tokens 2016, and about 300 tokens/s with batch size 1400 and input tokens 64.
Dovetail [202], Twinpilots [261] and Vellaisamy et al. [262] accommodate GPU-CPU and the hierarchical memory architecture (GPU and CPU memory) to maximize CPU and GPU utilization and fully hide the data loading time behind the CPU compute time. On a server with an NVIDIA A10G GPU with 24GB memory and an AMD EPYC 7R32 CPU with 256GB memory, Twinpilots [261] can achieve $\sim$5 tokens/s and 12 tokens/s for Llama-13B with batch size 1 and 8. PRIMA.CPP [263] proposes a piped-ring parallelism with prefetching and a scheduler to split model layers to heterogeneous devices. It can achieve 18.52 tokens/s and 22.73 tokens/s for Llama3-8B and Qwen2.5-7B on edge SoCs like Apple M1 ($\sim$14W) and Kirin 9000 ($\sim$7.9W). HeteroLLM [109] is a fast LLM inference engine with W4A16 quantization in mobile devices which supports both layer-level and tensor-level heterogeneous execution. Evaluation results show that HeteroLLM achieves 13-15 tokens/s for Llama2-7B on SnapDragon 8 Gen3 platform.
3.5.3 GPU
Memory Enhancement. (1) CPU-GPU. TightLLM [264] proposes and implements an adaptive offloading-based inference system to maximize inference throughput with two key innovations. Its key-value (KV) distributor employs a trade-compute-for-transfer strategy to address growing transfer overhead by dynamically recomputing portions of the KV cache and its weight loader slices model weights and distributes the loading process across multiple batches, effectively overlapping data transfer with computation and amortizing the excessive weight loading overhead. For the OPT-13B and Llama2-13B models, it achieves 22 and 20 tokens/s with maximum batch size on a NVIDIA RTX 3090 GPU and a 24-core CPU.
3.5.4 FPGA
Computing Enhancement. (1) FPGA-GPU. In latency-sensitive scenarios, a small batch or even one batch is usually required. This leads to the prefill and the decode stage of LLM inference being computational and memory bottlenecks, respectively. Therefore, it is difficult for a homogeneous FPGA or GPU system to simultaneously address different computational bottlenecks in different stages of LLM inference, resulting in long prefill latency on FPGAs and low utilization during the decode stage on GPUs. GLITCHES [265] is GPU-FPGA LLM inference through a collaborative heterogeneous system by employing GPUs for the prefill stage and FPGAs for the decode stage. Experiments demonstrate that a GLITCHES heterogeneous LLM inference system with an A100/V100S GPU and 7xU280 FPGAs achieves 390-450 tokens/s for Llama2-7B. HPU [266] proposes the High-Bandwidth Processing Unit (HPU) and GPU-HPU heterogeneous system, and develops a real FPGA-based HPU prototype with HBM.
3.5.5 PIM/NDP
@@
Computing Enhancement. (1) PIM-NPU. NeuPIMs [267] and IANUS [268] are heterogeneous PIM acceleration systems that jointly exploits a conventional GEMM-focused NPU and GEMV-optimized PIM devices. NeuPIMs [267] first proposes dual row buffers in each bank, facilitating the simultaneous management of memory read/write operations and PIM commands, to enable concurrent operations on both NPU and PIM platforms. Further, it employs a runtime sub-batch interleaving technique to maximize concurrent execution for the inherent dependencies between GEMM and GEMV in LLMs. Our evaluation demonstrates that NeuPIMs with $\sim$76W (0.6348W+75W) power consumption (32GB HBM) achieves $\sim$3k tokens/s with batch size 8 for GPT3-7B. IANUS [268] proposes novel PIM access scheduling that manages not only the scheduling of normal memory accesses and PIM computations but also workload mapping across the PIM and the NPU. The evaluations show that for GPT-6.7B model, IANUS achieves 127.1 tokens/s with about 240W power consumption. Cambricon-LLM [279] proposes a chiplet-based hybrid architecture with NPU and a dedicated NAND flash chip to enable efficient LLM inference. It utilizes both the high computing capability of NPU and the data capacity of the NAND flash chip. The NAND flash chip is enhanced by in-flash computing to perform precise lightweight on-die processing, and the NPU performs matrix operations and handles special function computations. Experimental results show that Cambricon-LLM with $\sim$37W power consumption achieves 3.44 tokens/s ($\sim$) and 36.34 tokens/s for 70B and 7B LLMs, which is over 22$\times$ to 45$\times$ faster than existing flash-offloading technologies, respectively. (2) PIM-GPU. MoNDE [269] is a near-data computing solution that efficiently enables Mixture-of-Experts (MoE) LLM inference on heterogeneous hardwares. MoNDE reduces the volume of MoE parameter movement by transferring only the hot experts to the GPU, while computing the remaining cold experts inside the host memory device. MoNDE can achieve inference latency comparable to an ideal GPU system with infinite memory. AttAcc [270, 271] is also an heterogeneous system equipped with GPUs to accelerate the attention layers. Compared to the monolithic state-of-the-art GPU system, AttAcc achieves significantly higher throughput (up to 2.81$\times$) and energy efficiency (up to 2.67$\times$) for GPT3-175B model. In 2024, SK Hynix proposes LPDDR6-based heterogeneous LLM system called AiMX-xPU[278], which consists 1 NVIDIA H100 GPU and 3 AiMX, achieving 167 tokens/s, 220 tokens/s, and 900 tokens/s with batch size 1, 8, and 32 for OPT-30B. PAISE [280] employs GPU-PIM heterogeneous computing resources to optimize LLM inference. For Llama2-7B model, the AMD MI100 GPU with HBM-PIM devices achives $\sim$400 tokens/s with batch size 1 and the power consumption is similar to the GPU-only system. Pyramid [282] optimizes GPU-PIM system for LLM inference by strategically allocating cross-level computational resources within PIM to meet diverse needs and leveraging the strengths of both technologies. Evaluation results demonstrate that Pyramid outperforms existing systems like NeuPIM, AiM, and AttAcc by factors of 2.31×, 1.91×, and1.72×, respectively. The inherent activation sparsity in LLMs naturally divides weight parameters into two categories, termed “hot” and “cold” neurons, respectively. Hot neurons consist of only $\sim$20of all weight parameters but account for 80% of the total computational load. In contrast, cold neurons account for just 20% of the computational workload. Hermes [283] leverages NDP units within DRAM DIMMs to enhance the performance of GPU by mapping hot neurons to GPU, while offloading cold neurons to NDP-DIMMs. Furthermore, Hermes introduce a lightweight predictor that ensures optimal real-time neuron partition and adjustment between GPU and NDP-DIMMs. Hermes facilitates the deployment of LLaMA2-70B on consumer-grade hardware at a rate of 13.75 tokens/s. (3) PIM-FPGA. Kang et al. [272] put the memory-bound GEMV calculations including projections, feed-forward-network layer, and the last fully-connected layer with pre-trained weight on HBM2-PIM and evaluate system with PIM-powered Xilinx Alveo U280 board. For GPT-1.3B model, after scaling the bandwidth by 8$\times$ for PIM technology, it can achieve about 347.83 tokens/s, which is 1.6$\times$ faster than the NVIDIA A100 GPU. From 2022 to 2023, SK Hynix proposes GDDR6-based accelerator-in-memory named AiM [303, 304] and integrates it with 2 Xilinx XCVU9P FPGA chips for control and communication as a prototype called AiMX [277] to achieve 330 tokens/s for OPT-6.7B with batch size 1. (4) PIM-TPU. H3D-Transformer [274] proposes a heterogeneous 3D-based accelerator design for transformer models, which adopts an interposer substrate with multiple 3D memory/logic hybrid cubes optimized for accelerating different MatMul workloads. An approximate computing scheme is proposed to take advantage of heterogeneous computing paradigms of mixed-signal compute-in-memory (CIM) and digital tensor processing units (TPU). From the system-level evaluation results, 10 TOPS/W energy efficiency is achieved for the GPT2 model, which is about 2.6$\times$-3.1$\times$ higher than the baseline with 7nm TPU and stacked FeFET memory. 3D-HI [276] leverage chiplet-based heterogeneous integration to design a Network-on-Interposer (NoI) architecture to accelerate LLM inference. 3D-HI uses the streaming multiprocessors along with the associated memory controllers (SM-MCs) for multi-head attention and ReRAM cores for the feed-forward network, which optimize both achievable energy efficiency and throughput. Further, vertical integration on top of a 2.5D interposer helps to enhance overall system performance and alleviates the issue of memory bottlenecks. Experimental results demonstrate that 3D-HI lowers the latency and energy consumption by up to 22.8$\times$ and 5.36$\times$ with respect to an equivalent state-of-the-art chiplet-based platform. PIM-LLM [168] leverages analog processing-in-memory (PIM) architectures and digital systolic arrays to accelerate 1-bit MatMul operations in projection layers and high-precision MatMul operations in attention, respectively. It can achieve 42.82 tokens/s and 170 tokens/J for Llama-7B with 128 sequence length.
@@
Figure 12: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with heterogeneous cooperation.